OpenStudy (anonymous):

limit x - > 0, (tan^2x)/x Find the limit, I need some help

OpenStudy (anonymous):

where do you think you shall begin

OpenStudy (anonymous):

use l'hospitals rule

OpenStudy (anonymous):

Well, I tried using that, But I usually get stuck at this part:

hartnn (hartnn):

thats correct, now separate one sin x from sin^2 x and bring it about x, to get sin x/x

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{\sin x}{x}=1\] someone correct me if im wrong

hartnn (hartnn):

u are right.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{\tan^2x}{x} \] \[\lim_{x \rightarrow 0} \frac{1}{x}*{\tan^2x} \] \[\lim_{x \rightarrow 0} \frac{1}{x}*\frac{\sin^2x}{\cos^2x} \] \[\lim_{x \rightarrow 0} \frac{\sin x}{x}*\frac{\sin x}{\cos^2x}=\lim_{x \rightarrow 0} \frac{\sin x}{x}*\lim_{x \rightarrow 0} \frac{\sin x}{\cos^2x}\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{\sin x}{x}=1\] please dont ask me to prove this, its just something you learn and will have to memorize all you really need to do is solve \[\lim_{x \rightarrow 0} \frac{\sin x}{\cos^2x}\]

OpenStudy (anonymous):

if you're still confused, remember that sin(0)=0 & cos(0)=1

OpenStudy (anonymous):

Okay, I got all that. So, It'll be 0/1? =0?

hartnn (hartnn):

yup

OpenStudy (anonymous):

Thank you guys for the help :D

OpenStudy (anonymous):

no prob